# From Tokens to Thoughts: How LLMs and Humans Trade Compression for Meaning

### v4

@@ A.10 Correlation between Human Typicality Judgments and LLM Internal Cluster Geometry
A.11 Typicality and Cosine Similarity [RQ2]
@@
Figure 16: Weak and Mostly Non-Significant Spearman Correlation Values Between Human Typicality Judgments and LLM Cosine Similarity Indicating Different Structure Representing Concepts. Mean Peak AMI Layer Spearman correlation values across the models belonging to the same family and across the three datasets.
A.12 Theoretical Extreme Case Exploration for $\mathcal{L}$
@@ A.13 Compression Figures
Figure 17: Static Embeddings Achieve a more “Optimal” Compression-Meaning Trade-off by the $\mathcal{L}$ Measure. IB-RDT objective ($\mathcal{L}$) vs. $K$ across all datasets. Lower $\mathcal{L}$ indicates a more optimal balance between compression ($I(X;C)$) and semantic fidelity (distortion). Static embeddings consistently achieve lower $\mathcal{L}$ values than both human categories and contextual embeddings. The plots correspond to the three datasets in the following order: Rosch [1973a], Rosch [1975], McCloskey and Glucksberg [1978].
Figure 18: Contextual Embeddings Achieve Better-than-Human Compression-Meaning Trade-off by the $\mathcal{L}$ Measure. IB-RDT objective ($\mathcal{L}$) vs. $K$ across all datasets. Lower $\mathcal{L}$ indicates a more optimal balance between compression ($I(X;C)$) and semantic fidelity (distortion). Contextual embeddings outperform human categories but achieve higher $\mathcal{L}$ values than static embeddings. The plots correspond to the three datasets in the following order: Rosch [1973a], Rosch [1975], McCloskey and Glucksberg [1978].
A.14 $\mathcal{L}$ objective vs. downstream task performance

### v5

@@ B.10 Correlation between Human Typicality Judgments and LLM Internal Cluster Geometry
B.11 Typicality and Cosine Similarity [RQ2]
@@
B.12 Theoretical Extreme Case Exploration for $\mathcal{L}$
@@ B.13 Compression Figures
Figure 17: Static Embeddings Achieve a more “Optimal” Compression-Meaning Trade-off by the $\mathcal{L}$ Measure. IB-RDT objective ($\mathcal{L}$) vs. $K$ across all datasets. Lower $\mathcal{L}$ indicates a more optimal balance between compression ($I(X;C)$) and semantic fidelity (distortion). Static embeddings consistently achieve lower $\mathcal{L}$ values than both human categories and contextual embeddings. The plots correspond to the three datasets in the following order: Rosch (1973a), Rosch (1975), McCloskey & Glucksberg (1978).
Figure 18: Contextual Embeddings Achieve Better-than-Human Compression-Meaning Trade-off by the $\mathcal{L}$ Measure. IB-RDT objective ($\mathcal{L}$) vs. $K$ across all datasets. Lower $\mathcal{L}$ indicates a more optimal balance between compression ($I(X;C)$) and semantic fidelity (distortion). Contextual embeddings outperform human categories but achieve higher $\mathcal{L}$ values than static embeddings. The plots correspond to the three datasets in the following order: Rosch (1973a), Rosch (1975), McCloskey & Glucksberg (1978).
B.14 $\mathcal{L}$ objective vs. downstream task performance